# Designing a compact cavity-enhanced source of entangled photons

### v1

@@ Abstract
Fig. 3(a) shows the SPDC bandwidth and the cluster spacing of the modes originating in crystal 1 for type-II and type-0 PM, as a function of the joint crystal length  [^2] Type-0 cluster spacing was calculated using a second order expansion, see Appendix C. For the SPDC photons originating in crystal 2, identical considerations can be made.. For type-0 SPDC, the bandwidth and the cluster spacing are exceedingly high, whereas for type-II SPDC, the SPDC bandwidth is notably less than the cluster spacing. This means that a source based on type-0 PM is not possible but a setup based on type-II may be feasible. In particular, using type-II PM can ensure that cavity enhancement is achieved for the intended single pair of frequencies with high fidelity.
@@
Secondly, let us consider the case of a highly non-degenerate source. We assume a pump wavelength of 519519519\,519nm to generate signal photons at the Rubidium transition $\lambda_{s}=780.24\,$nm and idler photons at $\lambda_{i}=1550\,$nm.
@@
Figure 4: Cluster spacings vs SPDC bandwidths for highly non-degenerate wavelengths $\lambda_{i}=780.24\,$nm and $\lambda_{s}=1550\,$nm. (a) Comparison of type-II and type-0 phase matching (PM) when creating photon pairs in a single crystal. The dashed lines show the bandwidths of the SPDC processes, the solid lines show the corresponding cluster spacings as a function of the joint crystal length $l_{1}+l_{2}$ at a fixed ratio $l_{1}/(l_{1}+l_{2})=0.4$. As in Fig. 3, the SPDC bandwidth is larger than the cluster spacing for type-0 PM. Therefore, type-II PM is more suitable for generating photon pairs of well-defined frequencies. (b) Comparing the joint cluster spacing to the individual cluster spacings and the SPDC bandwidths in the two crystals for type-II PM. We plot them as a function of the ratio of crystal lengths for a fixed joint crystal length $l_{1}+l_{2}=10\,$mm. In contrast to Fig. 3, there is a range of crystal length ratios (between $0.3$ and $0.4$), where the joint cluster spacing does not diverge but is still sufficiently large compared to the individual cluster spacings and SPDC bandwidths to create high-fidelity entanglement.
@@
Fig. 4(a) shows the SPDC bandwidth and the cluster spacing of a double resonance in crystal 1 as a function of the joint crystal length. Note that both the cluster spacing and the SPDC bandwidth reduce as the crystal length increases. For both types of phase matching, the SPDC bandwidth is too large to generate photon pairs where the frequency of each individual photon is well defined. The corresponding fidelity will therefore be only $0.5$ for a type-II source and $0.2$ for a type-0 source (see Fig. 2(b)). Type-0 PM has the additional disadvantage that the joint cluster spacing is on the order of THz (see Appendix B). This makes it impossible to find the quadruple resonances necessary for generating entangled photon pairs.
@@
We will therefore focus on finding quadruple resonances for the more favorable type-II PM. To achieve that, we investigate the dependence of the joint cluster spacing on the ratio of the crystal lengths as shown in Fig. 4(b). If the crystal lengths are the same, the joint cluster spacing will diverge since the two pairs of cavity modes become identical and will have the same cluster spacing for double resonances. To allow finding a quadruple resonance to generate entangled states, we therefore choose crystals of differing lengths. Even then, the SPDC processes in the individual crystals can still produce photons in multiple frequency modes.
@@
One solution is to introduce additional filtering of the cavity output. The practicality of this solution will depend on the spacing of the double resonances. Alternatively, one can introduce a relative optical path delay $\Delta l$ to idler modes $2$ and $4$ while keeping the signal modes unchanged. We model this path delay by adding $\Delta l\,\omega_{i}/\pi c$ to $m_{2/4}$. This will leave the SPDC bandwidths and the joint cluster spacing unchanged, but it will affect the cluster spacings for the SPDC processes in the individual crystals. Using this approach and $\Delta l=550\,\mu$m, one can achieve a fidelity of $0.93$ of the state generated with the targeted polarization entangled state as in Eq. 1 (see Appendix D).
The bandwidth of the state the source will generate depends on the cavity linewidth, which is determined by the intracavity losses and the cavity length. To provide a numeric estimate, let us assume the nonlinear crystals to be PPLN with a joint crystal length of 101010\,10mm. This will determine the optimal length of the cavity (535353\,53mm) in reference to the focusing parameter Boyd and Ashkin (1966). For mirror reflectivites of $99.98\%$, achieving a bandwidth of 111\,1MHz will require a finesse of $3000$. This gives an upper boundary of $0.2\%$ for the intracavity losses. Achieving such low losses is possible by picking appropriate anti-reflection coatings for PPLN. For additional details, see Appendix E.
We presented a design for a compact cavity-enhanced source of polarization-entangled photons. The photon pairs are generated in two nonlinear crystals with perpendicular optical axes inside a single Fabry-Pérot cavity. We studied the feasibility of the design for different types of phase matching and for near-degenerate as well as highly non-degenerate wavelengths. The combination of type-II phase matching and highly-nondegenerate wavelengths should allow the generation of high-fidelity narrowband entangled photon pairs for realistic experimental parameters. The presented design is the basis for an experimental realization we are currently working on. In addition to the considerations presented here, an actual implementation will require a high-precision control of experimental parameters to find and actively stay on the necessary quadruple resonance of the cavity. This will include the stabilization of the crystal temperatures, the cavity length, and the pump wavelength.
@@ Acknowledgements.
Appendix A Phase-mismatch and bandwidth
The bandwidth of a single-pass SPDC is determined by the phase mismatch of the nonlinear process $\Delta k=k_{s}+k_{i}-k_{p}-2\pi/\Lambda$. Here, $\Lambda$ is the periodic poling of the SPDC crystal. The amplitude of the photons generated in a crystal of length $l$ is proportional to $(\sin\xi/\xi)^{2}$ with $\xi=\Delta kl/2$. The FWHM bandwidth $\Delta\nu_{\text{SPDC}}$ of the SPDC photons can then be determined from $\xi_{\text{HWHM}}\approx 1.39$. Expanding $\Delta k$ up to linear order in signal frequency and temperature yields:
Here, $T$ is the crystal temperature, which affects the refractive index and the poling of the nonlinear crystal Jundt (1997); Gayer et al. (2008). The SPDC bandwidth is then determined from the solution of Eq. 5 as 2 $\omega_{s}$.
Appendix B Additional cluster spacing plots
Figure 5: (a) Cluster spacing as a function of the ratio of crystal lengths for a non-degenerate type-0 source at $\lambda_{i}=1550\,$nm and the Rubidium transition at $\lambda_{s}=780.24\,$nm. For all length ratios, the joint cluster spacing is larger than 111\,1THz, which makes the quadruple resonance essentially impossible to find. (b) Cluster spacing and SPDC bandwidth of a type-2 source at the Rubidium transition at $\lambda_{s}=780.24\,$nm and for $\lambda_{i}=1550\,$nm. An additional optical-path delay $\Delta l$ is added to modes 2 and 4. For $\Delta l=550\,\mu$m, the cluster spacings of the double resonances from both SPDC processes become twice as large as the corresponding SPDC bandwidths. For these values, one can achieve a fidelity of $0.93$ with the targeted maximally entangled state $|\psi^{-}\rangle$.
Appendix C Cluster spacing using higher-order approximations
In the main text, we used a linear approximation for calculating the cluster spacing. For most of the cases we considered there, this approximation yields the same results as a more detailed, higher-order calculation. The latter uses the definition of cluster spacing by Eckardt et al Eckardt et al. (1991). They found that the sum of two mode numbers $m=m_{1}+m_{2}$ will always be very close to an integer value if one is on a double resonance. This results from the fact that the sum of the signal and idler frequencies has to be equal to the pump frequency.
@@ Appendix C Cluster spacing using higher-order approximations
Apart from the frequencies of the optical modes involved, the cluster spacing can also depend on other system parameters. To take this into account, Eckert et al Eckardt et al. (1991) expand the joint mode number to second order in frequency and to first order in other parameters like, e.g., the crystal temperature $T$:
$\begin{split}\Delta m=&\left(\frac{\partial m}{\partial\omega_{j}}\right)_{\omega_{p}}\Omega+\\ &\frac{1}{2}\left(\frac{\partial^{2}m}{\partial\omega^{2}_{j}}\right)_{\omega_{p}}\Omega^{2}+\left(\frac{\partial m}{\partial T}\right)\left(T-T_{0}\right).\end{split}$ (7)
@@
The cluster spacing can be calculated by solving Eq. 6. This yields four possible solutions. We consider the solution yielding the lowest cluster spacing frequency because this determines the distance to the next closest cluster. In Fig. 6, we provide a comparison of the cluster spacing using the linear approximation of Eq. 3 in the main text and the cluster spacing derived from integer changes of the joint mode number. One can see deviations between the solutions for the first and second-order approximations close to parameters where the cluster spacing diverges. Examples are the case of a near-degenerate source close to equal crystal lengths ($l_{1}\approx l_{2}$) and the case of non-degenerate frequencies at $l_{1}\to 0$.
Appendix D The fidelity with the target Bell state
@@
Here, we assume that $\omega_{s}$ and $\omega_{i}$ both are resonant with the cavity. The amplitude for a given combination of signal and idler frequencies to the overall state is given by the normalized amplitude $A_{j}^{\alpha}=\mathrm{sinc}(l_{\alpha}\Delta k_{j,\alpha}/2)$ of the SPDC process at that frequency. The index $\alpha=1,2$ denotes whether the SPDC process is happening in the first or the second crystal. If there is only one crystal as in the example in the main text, one can drop the index $\alpha$ and write $\Delta k_{j}$ instead. To first order, the phase mismatch is equal to:
The pump polarization is chosen such that the absolute values of the amplitudes for the central frequencies are equal ($|A_{0}^{1}|=|A_{0}^{2}|$), and that their relative phase matches that of the target Bell state. For example, $A_{0}^{1}=-A_{0}^{2}$ for $|\psi^{-}\rangle$. All of the other frequency contributions with $n\neq 0$ lower the fidelity with the targeted Bell state, $F=|\langle\Psi|\psi^{\pm}\rangle|^{2}$.
Appendix E The minimum linewidth in a cavity
The goal of the present paper has been to describe the design of a source of entangled photons allowing efficient coupling to quantum memories. This coupling efficiency will be better the narrower the photon bandwidth. However, that bandwidth has to balance two contrary needs: a high rate of entangled pairs to overcome losses and a good coupling efficiency to quantum memories. Here, we assume a bandwidth of 111\,1MHz to be a reasonable compromise. For example, this will allow good coupling to Rb quantum memories Prakash et al. (2019); Wang et al. (2019).
Figure 7: The finesse required to achieve a specific cavity linewidth as a function of the cavity length.
The bandwidth achievable for the photon pairs produced will be determined by the properties of the cavity. In particular, it is related to the finesse, which is defined as:
$\mathcal{F}=\frac{\textrm{FSR}}{\Delta\nu},$ (12)
where $\Delta\nu$ is the FWHM linewidth of the cavity, and $\textrm{FSR}=c/2L_{\textrm{cavity}}$. The dependence of the finesse on the cavity length $L_{\textrm{cavity}}$ and its relation to the cavity linewidth are shown in Fig. 7. Of course, increasing the cavity length too much will contradict our goal of building a compact source of entangled photons.
To keep the source as compact as possible, one has to find other means of increasing the cavity finesse. The finesse depends on the mirror reflectivities $R_{1}$ and $R_{2}$ and the intracavity losses $\eta$ in the following way:
A derivation of an equivalent expression for the finesse can be found in Ref. Ismail et al. (2016). Increasing the finesse can be achieved by choosing high reflectivities for the cavity mirrors, high-quality antireflection coatings for the crystal end faces and low-absorption nonlinear crystals.
Let us now estimate an upper limit on the reflectivities of the antireflection coating for the wavelengths of the SPDC photons. To this end, we fix the two mirror reflectivities to be the same ($R_{1}=R_{2}\equiv R$) and then calculate the linewidth as a function of the corresponding intra-cavity losses using Eq. (12). &  (13). In Fig. 8, this dependence is shown for a 111\,1cm long PPLN crystal and for several choices of mirror reflectivities. We neglected the losses in PPLN compared to the losses due to the anti-reflection coatings. The cavity length is determined by first finding the optimal focusing condition Boyd and Kleinman (1968) and then matching the curvature of the optimally focused beam to the curvature of the cavity mirrors, which we assume to be 151515\,15mm. By also taking into account the refractive index of PPLN, we get an effective cavity length of 535353\,53mm.
Figure 8: The dependence of the cavity linewidth on the intracavity losses for specific reflectivities of the cavity mirrors. The cavity length is 535353\,53mm.

### v2

@@ IV.1 Near-degenerate source
Fig. 3(a) shows the SPDC bandwidth and the cluster spacing of the modes originating in crystal 1 for type-II and type-0 PM, as a function of the joint crystal length  [^2] Type-0 cluster spacing was calculated using a second order expansion. See Supplemental Material [58] and  [59] for more details. For the SPDC photons originating in crystal 2, identical considerations can be made.. For type-0 SPDC, the bandwidth and the cluster spacing are exceedingly high, whereas for type-II SPDC, the SPDC bandwidth is notably less than the cluster spacing. This means that a source based on type-0 PM is not possible but a setup based on type-II may be feasible. In particular, using type-II PM can ensure that cavity enhancement is achieved for the intended single pair of frequencies with high fidelity.
@@ IV.2 Highly non-degenerate source
Secondly, let us consider the case of a highly non-degenerate source. We assume a pump wavelength of 519519519\,519nm to generate signal photons at the Rubidium transition $\lambda_{s}=780.24\,$nm and idler photons at $\lambda_{i}=1550\,$nm. The idler photons would then be suitable for a long-distance distribution over telecom fibers. If they are sufficiently narrowband, one could also couple the idler photons to optomechanical systems.
@@
Figure 4: Cluster spacings vs SPDC bandwidths for highly non-degenerate wavelengths $\lambda_{i}=780.24\,$nm and $\lambda_{s}=1550\,$nm. (a) Comparison of type-II and type-0 phase matching (PM) when creating photon pairs in a single crystal. The dashed lines show the bandwidths of the SPDC processes, the solid lines show the corresponding cluster spacings as a function of the joint crystal length $l_{1}+l_{2}$ at a fixed ratio $l_{1}/(l_{1}+l_{2})=0.4$. (b) Comparing the joint cluster spacing to the cluster spacings and the SPDC bandwidths in individual crystals for type-II PM. We plot them as a function of the ratio of crystal lengths for a fixed joint crystal length $l_{1}+l_{2}=10\,$mm. (c) Comparing the joint cluster spacing to the cluster spacings and the SPDC bandwidths in individual crystals for type-0 PM. For all length ratios, the joint cluster spacing is larger than 111\,1THz, which makes the quadruple resonance essentially impossible to find. (d) Cluster spacing and SPDC bandwidth of a type-II source. An additional optical-path delay $\Delta l$ is added to modes 2 and 4.
@@
Fig. 4(a) shows the SPDC bandwidth and the cluster spacing of a double resonance in crystal 1 as a function of the joint crystal length. Note that both the cluster spacing and the SPDC bandwidth reduce as the crystal length increases. For both types of phase matching, the SPDC bandwidth is too large to generate photon pairs where the frequency of each individual photon is well defined. The corresponding fidelity will therefore be only $0.5$ for a type-II source and $0.2$ for a type-0 source (see Fig. 2(b)). Type-0 PM has the additional disadvantage that the joint cluster spacing is on the order of THz, as shown in Fig. 4(c). This makes it impossible to find the quadruple resonances necessary for generating entangled photon pairs.
@@
We will therefore focus on the more favorable type-II PM. To achieve that, we investigate the dependence of the joint cluster spacing on the ratio of the crystal lengths as shown in Fig. 4(b). If the crystal lengths are the same, the joint cluster spacing will diverge since the two pairs of cavity modes become identical and will have the same cluster spacing for double resonances. To allow finding a quadruple resonance to generate entangled states, we therefore choose crystals of differing lengths. Fig. 4(b) shows that a length ratio between 0.3 and 0.4 provides reasonable values for the joint cluster spacing. In this range, it is still sufficiently small to allow finding a quadruple resonance for realistic experimental parameters. Even then, the SPDC processes in the individual crystals can still produce photons in multiple frequency modes.
@@
We suggest two solutions to ensure a high quality of the entanglement generated. One is to use a compact cavity design in combination with moderate spectral filtering. The practicality of this solution will depend on the spacing of the double resonances. The second solution is to introduce a relative optical path delay between the signal and idler modes to increase the individual cluster spacings. While this will result in a more complex and larger cavity design, it will allow generating narrowband entangled photons without the need for any additional spectral filtering. If the complexity and scale of the source do not pose a limitation, larger cavity lengths can allow even significantly narrower linewidths. In particular, we introduce an optical path delay $\Delta l_{2}$ to idler mode $2$ and $\Delta l_{4}$ to mode $4$. We model this by adding the term $\Delta l_{2/4}\,\omega_{i}/\pi c$ to the mode numbers $m_{2/4}$. This will leave the SPDC bandwidths and the joint cluster spacing unchanged, but it will affect the cluster spacings for the SPDC processes in the individual crystals. The comparison of cluster spacings and SPDC bandwidths as a function of path delay is shown in Fig. 4(d). For simplicity, we take $\Delta l_{2}=\Delta l_{4}=\Delta l$. For $\Delta l=550\,\mu$m, the cluster spacings of the double resonances from both SPDC processes become twice as large as the corresponding SPDC bandwidths. For these values, one can achieve a fidelity of $0.93$ with the targeted maximally polarization entangled state $|\psi^{-}\rangle$.
V Cavity parameters
The bandwidth of the photons leaving the cavity is determined by the linewidth of the cavity mode, which in turn will depend the intracavity losses and the cavity length. These are the parameters one can use to optimize the cavity characteristics and define the bandwidth of the photons generated. In the following, we will discuss bandwidths achievable using realistic experimental parameters. For our calculation, we assume the nonlinear crystals to be PPLN with a joint crystal length of 101010\,10mm.
Figure 5: The finesse required to achieve a specific cavity linewidth as a function of the cavity length. In practice the desire for increased cavity length needs to be balanced with a need for a stable and compact cavity design.
The FWHM linewidth $\Delta\nu$ of the cavity is related to the finesse, which is defined as:
$\mathcal{F}=\frac{\textrm{FSR}}{\Delta\nu},$ (8)
where $\textrm{FSR}=c/2L_{\textrm{cavity}}$ for a standing-wave resonator. The dependence of the finesse on the cavity length $L_{\textrm{cavity}}$ and its relation to the cavity linewidth are shown in Fig. 5. While increasing the cavity length will allow narrower linewidths, this will negatively impact the stability and compactness of our source of entangled photons.
If we assume a simple Fabry-Pérot cavity with two mirrors of an identical radius of curvature, the optimal cavity length can be determined as follows. One first finds the optimal focusing condition [52, 53] and then matches the curvature of the optimally focused beam to the curvatures of the cavity mirrors. If one assumes the radii of curvature to be 151515\,15mm, and if one takes into account the refractive index of PPLN, one arrives at an effective cavity length of 535353\,53mm.
Once the cavity length is determined, one can aim to decrease the cavity linewidth by choosing appropriate mirror reflectivities and by minimizing the intra-cavity losses. In particular, the finesse depends on the mirror reflectivities $R_{1}$ and $R_{2}$ and the intracavity losses $\eta$ in the following way:
A derivation of an equivalent expression for the finesse can be found in Ref. [55]. Increasing the finesse can be achieved by decreasing the overall losses in the cavity through choosing high reflectivities for the cavity mirrors, high-quality antireflection coatings for the crystal end faces and low-absorption nonlinear crystals.
Let us now estimate an upper limit on the reflectivities of the antireflection coating for the wavelengths of the SPDC photons. To this end, we fix the two mirror reflectivities to be the same ($R_{1}=R_{2}\equiv R$) and then calculate the linewidth as a function of the corresponding intra-cavity losses using Eq. (8). and Eq. (9). In Fig. 6, this dependence is shown for a 111\,1cm long PPLN crystal and for several choices of mirror reflectivities. We neglected the losses in PPLN (less than 0.0005 $\mathrm{cm}^{-1}$ for considered wavelengths [56]) compared to the losses due to the anti-reflection (AR) coatings (for example, crystals with 0.2% AR coating are available from manufacturer HC Photonics).
Figure 6: The dependence of the cavity linewidth on the intracavity losses for specific reflectivities of the cavity mirrors. The cavity length is 535353\,53mm. By picking the desired cavity linewidth and the available mirror reflectivities one can determine the criteria for allowed losses in the cavity. Manufacturers usually only provide upper boundaries for reflectivities and antireflection coatings, which needs to be acknowledged when designing the cavity.
For example, if we assume $0.2\%$ AR coatings, the total intracavity losses will be about $0.8\%$ because we of the four air-to-crystal boundaries the photons will encounter per round-trip. This will result in a cavity linewidth of about 444\,4MHz for mirror reflectivities of $99.98\%$.
VI Finding and stabilizing the cavity resonances
Maintaining resonances of the cavity requires active stabilization of the cavity. This can be done using a standard Pound-Drever-Hall technic [57]. Because the four resonances are not independent due to energy conservation of SPDC, it will be sufficient to prevent the drift of two modes to stabilize the overall system. Stabilizing one mode prevents an absolute drift of the set of resonances, while stabilizing the second mode prevents relative drifts. The piezo on one of the cavity mirrors can follow the cavity drifts to maintain it on resonance with the reference pump light. The reference laser can be locked to the second mode to prevent the relative positions of resonances drifting apart.
Locking can be done in two regimes. One option is to use a continuous lock. In this case, the cavity resonances are stabilized in reference to error signals of resonances that are one FSR detuned from the operating wavelengths. These strong laser signals can then be filtered from the single-photon modes using filtering cavities without affecting the entangled state. The suppression of light from the locking lasers can be enhanced by using a bow-tie cavity design, which allows an additional spatial distinction between counter-propagating modes [18]. Another way to lock is to operate the cavity alternatingly in two regimes [23]: (1) an optical parametric oscillator (OPO) regime for locking, where strong reference laser light above the OPO threshold pumps the crystals inside the cavity. (2) Periods of time during which entangled pairs are generated. In regime (2), no reference light is present, and the cavity is left to drift freely for a short time. This requires appropriately fast optical switches, intensity modulators and circulators to protect single photon detectors from strong reference laser light, and to switch the pump beam above and below the OPO threshold.
The circulators will also play an essential role in allowing the efficient collection of single photons generated inside the cavity into single mode fibers. This coupling to a single mode will act as a spatial filter for light coupled into and out of the cavity. The amount of light collected will primarily depend on the proper impedance matching of the cavity. This will also strongly depend on the choice of using a reflective or a transmissive cavity design.
As presented in previous sections, an important challenge will be to find a suitable quadruple resonance. Finding and stabilizing the cavity resonances can be achieved by controlling several degrees of freedom: 1) the temperatures of the nonlinear crystals 2) the frequency of the pump light, and 3) moving a cavity mirror with a piezo. For example, tuneable narrowline lasers typically offer several 10 GHz of mode-hop free frequency tuning. If one assumes a PPLN SPDC crystal of about 101010\,10mm length, changing its temperature by 111\,1K will shift the corresponding cluster spacing for the signal and idler frequencies by about 222\,2THz. This large dependence on temperature means that the crystal temperature has to be stabilized with mK precision. Tuning the temperature will be essential for initially finding a quadruple resonance. Active stabilization will be done using a piezo actuator glued to a cavity mirror. In this case, one has to take into account the piezo response curve and how it responds to small voltage fluctuations. The maximum displacement of the piezo has to allow shifting the cavity resonance by at least one FSR for the cavity modes one needs to stabilize. This is necessary to allow the cavity lock to compensate drifts at high and intermediate frequencies.
VII Conclusion
Our work presents a novel design for cavity-enhanced sources of polarization-entangled photons, where the photons are created in two SPDC crystals inside a single cavity. We provide a comprehensive study of the feasibility of the design for different types of phase matching and for the cases of near-degenerate as well as highly non-degenerate wavelengths. We find that combining type-II phase matching and highly-nondegenerate wavelengths should allow the generation of high-fidelity narrowband entangled photon pairs for realistic experimental parameters. Our approach allows achieving narrow bandwidths with a simple Fabry-Pérot cavity and only moderate spectral filtering to suppress the emission of unwanted frequency modes. No spectral filtering is needed if one adds dispersive elements to the cavity. While this results in a less compact cavity design, a longer cavity offers the opportunity to achieve a significantly narrower linewidth. Our approach provides essential considerations when implementing narrow-band cavity-enhanced SPDC sources of entanglement for a variety of applications in future quantum networks.
@@ Acknowledgements.
Appendix A Cluster spacing using higher-order approximations
In the main text, we used a linear approximation for calculating the cluster spacing. For most of the cases we considered there, this approximation yields the same results as a more detailed, higher-order calculation. The latter uses the definition of cluster spacing by Eckardt et al [59]. They found that the sum of two mode numbers $m=m_{1}+m_{2}$ will always be very close to an integer value if one is on a double resonance. This results from the fact that the sum of the signal and idler frequencies has to be equal to the pump frequency.
@@ Appendix A Cluster spacing using higher-order approximations
Apart from the frequencies of the optical modes involved, the cluster spacing can also depend on other system parameters. To take this into account, Eckert et al [59] expand the joint mode number to second order in frequency and to first order in other parameters like, e.g., the crystal temperature $T$:
$\Delta m=\left(\frac{\partial m}{\partial\omega_{j}}\right)_{\omega_{p}}\Omega+\frac{1}{2}\left(\frac{\partial^{2}m}{\partial\omega^{2}_{j}}\right)_{\omega_{p}}\Omega^{2}+\left(\frac{\partial m}{\partial T}\right)\left(T-T_{0}\right).$ (11)
@@
The cluster spacing can be calculated by solving Eq. 10. This yields four possible solutions. We consider the solution yielding the lowest cluster spacing frequency because this determines the distance to the next closest cluster. In Fig. 7, we provide a comparison of the cluster spacing using the linear approximation we used in the main text and the cluster spacing derived from integer changes of the joint mode number. One can see deviations between the solutions for the first and second-order approximations close to parameters where the cluster spacing diverges. Examples are the case of a near-degenerate source close to equal crystal lengths ($l_{1}\approx l_{2}$) and the case of non-degenerate frequencies at $l_{1}\to 0$.
Appendix B The fidelity with the target Bell state
@@
Here, we assume that $\omega_{s}$ and $\omega_{i}$ both are resonant with the cavity. The amplitude for a given combination of signal and idler frequencies to the overall state is given by the normalized amplitude $A_{j}^{\alpha}=\mathrm{sinc}(l_{\alpha}\Delta k_{j,\alpha}/2)$ of the SPDC process at that frequency. The index $\alpha=1,2$ denotes whether the SPDC process is happening in the first or the second crystal. The pump polarization is chosen such that the absolute values of the amplitudes for the central frequencies are equal ($|A_{0}^{1}|=|A_{0}^{2}|$), and that their relative phase matches that of the target Bell state. For example, $A_{0}^{1}=-A_{0}^{2}$ for $|\psi^{-}\rangle$. All of the other frequency contributions with $n\neq 0$ lower the fidelity with the targeted Bell state, $F=|\langle\Psi|\psi^{\pm}\rangle|^{2}$.